OpenStudy (anonymous):

ordered pairs for the linear equation 2x+5y=3

geerky42 (geerky42):

Exactly what do you need? There is infinitely ordered pairs

OpenStudy (precal):

just pick any three pairs

OpenStudy (anonymous):

you can easily find ordered pairs by plugging in values of your choice for either x or y and then solving for the other value. For example: \[2x+5y=3\] say \[x=2\] so \[2(2)+5y=3\] \[4+5y=3\] \[5y=-1\] \[y=-1/5\] so the ordered pair is (2, -1/5). You can pick whatever values best fit your needs.